mikewwe13:

The table of values represents a quadratic function. What is the the average rate of change for f(x) from x= −5 to x = 10 ? Enter your answer in the box. _______

mikewwe13:

@Vocaloid

Vocaloid:

look at the table find the value f(10), the value f(-5) then f(10) - f(-5) = ?

mikewwe13:

1 attachment
mikewwe13:

you seen the pic ?

Vocaloid:

yes

mikewwe13:

15

Vocaloid:

remember what I said before, f(10) doesn't mean multiply f by 10, it means you need to find the value of the function at x = 10

mikewwe13:

204

Vocaloid:

good, now repeat the process for f(-5)

mikewwe13:

39

Vocaloid:

good, now, using *** the values of f(10) and f(-5) you just found *** then f(10) - f(-5) = ?

mikewwe13:

165

Vocaloid:

good, then we divide that by (10 - (-5)) so 165/ (10 - (-5)) = ?

mikewwe13:

11

Vocaloid:

good, so 11 = your answer

mikewwe13:

Let f(x) = −1/4(x + 4)^2 − 8 . What is the average rate of change for the quadratic function from x = −2 to x = 2? Enter your answer in the box. ________

Vocaloid:

same thing as before, f(2) - f(-2) *** the answer is not 4 or 4f, you need to find the values of f(2) and f(-2) individually by plugging x = 2 and x = -2 into the function.

mikewwe13:

-9 and -17

Vocaloid:

good, so f(2) - f(-2) = ?

mikewwe13:

the answer is -2

Vocaloid:

-17 - (-9) = ?

mikewwe13:

-8

Vocaloid:

good then divide that by (2 - (-2))

mikewwe13:

4

Vocaloid:

sorry

Vocaloid:

yeah it should be -4 not 4 since -8/2 = -4 = your answer

mikewwe13:

Which function grows at the fastest rate for increasing values of x? h(x) = 6x^2 + 5x + 20 f(x) = 5 ⋅ 3x g(x) = 50x + 35

Vocaloid:

is the x in f(x) an exponent?

Vocaloid:

like, is it 3^x or 3x?

mikewwe13:

i mean yes

Vocaloid:

ok, since the x is in the exponent, that means it will grow faster than just having an exponent of 2 or 1 like the others so f(x) is the fastest

mikewwe13:

cause when you multiply variables with exponents, you add the exponents

mikewwe13:

https://static.k12.com/nextgen_media/assets/8090372-NG_AL1_O_04_U12_Quiz_01.png The graph shows the quadratic function f(x) . What is the average rate of change for the quadratic function from x = 6 to x = 8? Enter your answer in the box. _______

Vocaloid:

I think we already did this one, we got -1 as our answer

mikewwe13:

we did ?

Vocaloid:

yes

Vocaloid:

I don't feel like looking for it but if you look through your last question it's in there somewhere